OpenStudy (anonymous):

What is the GCF of 40xy and 24x? 8x 4xy 24x 40xy

OpenStudy (aroub):

8x

OpenStudy (anonymous):

aroooub!

OpenStudy (anonymous):

40xy=2*2*2*5*x*y 24x=2*2*2*3*x so GCF=2*2*2*x =8x

OpenStudy (aroub):

vveerroonniiccaaa!!! <3

OpenStudy (aroub):

thanks for the medals..!!!! :D :D

OpenStudy (anonymous):

friend me on fb! and message me! <3

OpenStudy (anonymous):

the first one

OpenStudy (anonymous):

aroub add me too i just reactivated

OpenStudy (aroub):

hehehe finalllyyyyyyy!!! okaayyy

OpenStudy (aroub):

ok hahd whats ur full name??

OpenStudy (anonymous):

hahd khan

OpenStudy (anonymous):

try searching with email hahd@live.ca it usually works

OpenStudy (aroub):

oh ur khan too..!!! okay just a min! :)

OpenStudy (aroub):

nah.. didnt find you!! :( its says fahd and for the email hand@live.ca

OpenStudy (anonymous):

its hahd@live.ca tell me urs, fb put age restriction search in canada if u are under 18 then u cant be looked up unless u are a friend of a friend

OpenStudy (aroub):

still..! ok mine is aroub husari i guess im the only one which has this name...

OpenStudy (aroub):

oh and veronica i just did!! :)

OpenStudy (anonymous):

did u get mine cause like i was just wondering did i add the right person ?

OpenStudy (aroub):

yes!! :)